# Black-hole hair from vector dark matter accretion

### v1

@@ VI.2 Mass of cloud
$M_{c}(r)=\oint_{S^{2}}\textrm{d}\Omega\int_{r_{\rm{s}}}^{r}r^{\prime 2}\textrm{d}r^{\prime}\rho(r^{\prime},\theta,\phi)\,,$ (70)
which we interpret as the mass of dark matter enclosed between the horizon and the surface at $r$. This quantity is plotted for a range of field masses with a BH mass of $10^{9}M_{\odot}$ in Fig. 5. This quantity scales with the BH mass $M$ by $M_{c}\sim M^{3}$. Thus, to convert the numbers in Fig. 5 to those for a BH with a mass $M$, one should multiply them by $\left(M/10^{9}M_{\odot}\right)^{3}$. As discussed previously, we choose the ambient dark matter density to be $\rho_{c}\approx 10M_{\odot}/\text{pc}^{3}$, modeling a dark matter soliton in a galactic center.
@@ VI.3 Mass accretion rate
where $k^{\mu}\partial_{\mu}=f^{1/2}\partial_{r}$ and $u^{\mu}\partial_{\mu}=f^{-1/2}\partial_{t}$. This quantity is plotted for a range of $\mu M$ values for a BH of mass $M=10^{9}M_{\odot}$ in Fig. 7.
@@ VII.2 Contribution to black hole growth
Another astrophysical effect we highlight is the contribution of vector dark matter accretion to BH growth. As shown in Section VI.3, for the particle regime ($\mu M\gtrsim 1$), the central BH accretes mass at a rate dependent only on its mass $M$ and the ambient dark matter density $\rho_{c}$, given by Eq. (76). For a $M\sim 10^{9}M_{\odot}$ supermassive BH, this regime is reached for any field mass $\mu\gtrsim 10^{-19}$ eV—practically the entire mass range of interest. Moreover, some cosmological models, such as the inflationary production mechanism of Ref. [5], prefer vector dark matter masses roughly in the range $10^{-10}\text{ eV}<\mu\lesssim 10^{-4}\text{ eV}$. In this scenario, even stellar-mass BHs would sit in the particle regime, meaning every known BH in the universe would accrete vector dark matter at the rate given by Eq. (76).
@@ A.1 General $r\to\infty$ case
$\left(\frac{d^{2}}{dr^{2}}+\frac{r^{4}k^{2}+r^{3}\mu^{2}-r(r-1)\ell(\ell+1)+r-3/4}{r^{2}(r-1)^{2}}\right)y=0\,.$
@@
$\left(\frac{d^{2}}{dr^{2}}+k^{2}+\frac{2k^{2}+\mu^{2}}{r}+\frac{3k^{2}-\ell(\ell+1)+2\mu^{2}}{r^{2}}\right)y=0\,.$ (78)
@@ A.1.1 $r\to\infty,k=0$ case
$\left(\frac{d^{2}}{dr^{2}}+\frac{\mu^{2}}{r}+\frac{2\mu^{2}-\ell(\ell+1)}{r^{2}}\right)y=0\,.$ (84)
@@ B.2 $r\to\infty,k=0$ case
$\biggl{(}\frac{d^{2}}{dr^{2}}+\frac{\omega^{2}}{f^{2}}+\frac{1}{4r^{4}f^{2}}+\frac{3\nu^{2}}{q_{r}^{2}}+\frac{\nu^{2}}{rfq_{r}}{}$ (93)
@@ Appendix C Cloud mass and accretion rate
We now translate from a differential conservation law to an integral one. To do so, we introduce four 3-surfaces: the horizon surface =$\mathcal{H}=\{\mathcal{M}:r=r_{\rm{s}}\}$, a timelike surface $\chi=\{\mathcal{M}:r=r_{0}\}$, where $r_{0}$ is an arbitrary radius outside the horizon, and finally two spacelike surfaces at $t=t_{1},t_{2}$, $\Sigma_{i}=\{\mathcal{M}:t=t_{i}\}$. We call the spacetime region enclosed between these surfaces $\mathcal{V}$. Integrating the conservation law over the bulk and applying Stokes’ theorem, we find
[…]
where $d\Sigma_{\mu}$ is the directed area element of the boundary hypersurface. It is given by $d\Sigma_{\mu}=d^{3}x\sqrt{|\gamma|}n_{\mu}$, where $\gamma$ is the determinant of the induced metric, $n_{\mu}$ is the normal, and the overall sign is chosen based on the geometry of each segment. We may split $\partial\mathcal{V}$ into the surfaces $\partial\mathcal{V}=\tilde{\Sigma}_{1}\cup\tilde{\mathcal{H}}\cup\tilde{\Sigma}_{2}\cup\tilde{\chi}$, where the tilde denotes the segment of each surface intersecting $\partial\mathcal{V}$, or $\tilde{\Sigma}=\Sigma\cap\partial\mathcal{V}$ for generic surface $\Sigma$. We may thus rewrite the boundary integral equation (100) as
[…]
@@
[…]
@@
[…]
@@ Appendix D Self-consistency of decoupling approximation
The perturbative expansion applies when the Proca test field is “small” compared to the gravitational background. To define a notion of “smallness” we turn to the Einstein’s equations
@@
ϵitalic-$\epsilon\approx{}$ $\frac{r^{3}}{M}\rho$
[…]

### v2

@@ VI.2 Mass of cloud
$M_{c}(r)=\oint_{S^{2}}\textrm{d}\Omega\int_{r_{\rm{s}}}^{r}\textrm{d}r^{\prime}r^{\prime 2}\rho(r^{\prime},\theta,\phi)\,$, (71)
which we interpret as the mass of dark matter enclosed between the horizon and the surface at $r$. This quantity scales with the BH mass $M$ by $M_{c}\sim M^{3}$, so it should be most relevant for the largest BHs. We thus plot it for a $10^{9}M_{\odot}$ supermassive BH over a range of field masses in Fig. 5. To convert the numbers in Fig. 5 to those for a BH with a mass $M$, one should multiply them by $\left(M/10^{9}M_{\odot}\right)^{3}$. As discussed previously, we choose the ambient dark matter density to be $\rho_{c}\approx 10M_{\odot}/\text{pc}^{3}$, modeling a dark matter soliton in a galactic center.
@@ VI.3 Mass accretion rate
where $k^{\mu}\partial_{\mu}=f^{1/2}\partial_{r}$ and $u^{\mu}\partial_{\mu}=f^{-1/2}\partial_{t}$. This quantity is plotted for a range of $\mu M$ values for a BH of mass $M=10^{9}M_{\odot}$ in Fig. 7. Like with the cloud mass, this BH mass is chosen because the accretion rate is maximized for the largest BHs.
@@
We additionally see, as in Fig. 3, that angular momentum effects are irrelevant in the particle regime but become dominant in the wave regime. The curves plotted in Fig. 7 are the monopole ($\ell=0$) mode and the $\ell=1$ modes of the three polarizations. The vector-type E-mode has total angular momentum $j=\ell+S=0$, which is the same as the monopole mode. This gives the two modes nearly identical behavior in the wave regime. The B-mode and the scalar-type E-mode have total angular momenta of $j=1$ and $j=2$, respectively, and their accretion rates are exponentially suppressed in the wave regime. We thus see that in the wave regime, accretion is most relevant for the modes of lowest angular momentum, and it is suppressed for higher-$j$ modes.
@@ VII.2 Contribution to black hole growth
[…]
@@ A.1 General $r\to\infty$ case
$\left(\frac{\textrm{d}^{2}}{\textrm{d}r^{2}}+\frac{r^{4}k^{2}+r^{3}\mu^{2}-r(r-1)\ell(\ell+1)+r-3/4}{r^{2}(r-1)^{2}}\right)y=0\,.$
@@
$\left(\frac{\textrm{d}^{2}}{\textrm{d}r^{2}}+k^{2}+\frac{2k^{2}+\mu^{2}}{r}+\frac{3k^{2}-\ell(\ell+1)+2\mu^{2}}{r^{2}}\right)y=0\,.$ (79)
@@ A.1.1 $r\to\infty,k=0$ case
$\left(\frac{\textrm{d}^{2}}{\textrm{d}r^{2}}+\frac{\mu^{2}}{r}+\frac{2\mu^{2}-\ell(\ell+1)}{r^{2}}\right)y=0\,.$ (85)
@@ B.2 $r\to\infty,k=0$ case
$\biggl{(}\frac{\textrm{d}^{2}}{\textrm{d}r^{2}}+\frac{\omega^{2}}{f^{2}}+\frac{1}{4r^{4}f^{2}}+\frac{3\nu^{2}}{q_{r}^{2}}+\frac{\nu^{2}}{rfq_{r}}{}$ (94)
@@ Appendix C Cloud mass and accretion rate
We now translate from a differential conservation law to an integral one. To do so, we introduce four 3-surfaces: the horizon surface $\mathcal{H}=\{\mathcal{M}:r=r_{\rm{s}}\}$, a timelike surface $\chi=\{\mathcal{M}:r=r_{0}\}$, where $r_{0}$ is an arbitrary radius outside the horizon, and finally two spacelike surfaces at $t=t_{1},t_{2}$, $\Sigma_{i}=\{\mathcal{M}:t=t_{i}\}$. We call the spacetime region enclosed between these surfaces $\mathcal{V}$. Integrating the conservation law over the bulk and applying Stokes’ theorem, we find
$0={}$ $\int_{\mathcal{V}}\textrm{d}^{4}x\sqrt{-g}\nabla_{\mu}J^{\mu}=\oint_{\partial\mathcal{V}}\textrm{d}\Sigma_{\mu}J^{\mu}\,$, (101)
where $\textrm{d}\Sigma_{\mu}$ is the directed area element of the boundary hypersurface. It is given by $\textrm{d}\Sigma_{\mu}=\textrm{d}^{3}x\sqrt{|\gamma|}n_{\mu}$, where $\gamma$ is the determinant of the induced metric, $n_{\mu}$ is the normal, and the overall sign is chosen based on the geometry of each segment. We may split $\partial\mathcal{V}$ into the surfaces $\partial\mathcal{V}=\tilde{\Sigma}_{1}\cup\tilde{\mathcal{H}}\cup\tilde{\Sigma}_{2}\cup\tilde{\chi}$, where the tilde denotes the segment of each surface intersecting $\partial\mathcal{V}$, or $\tilde{\Sigma}=\Sigma\cap\partial\mathcal{V}$ for generic surface $\Sigma$. We may thus rewrite the boundary integral equation (101) as
$0={}$ $\int_{\tilde{\Sigma}_{2}}\textrm{d}\Sigma_{\mu}J^{\mu}-\int_{\tilde{\Sigma}_{1}}\textrm{d}\Sigma_{\mu}J^{\mu}$
$+\int_{\tilde{\mathcal{H}}}\textrm{d}\Sigma_{\mu}J^{\mu}-\int_{\tilde{\chi}}\textrm{d}\Sigma_{\mu}J^{\mu}\,$, (102)
@@
$\mathcal{E}_{i}={}$ $\int_{\tilde{\Sigma}_{i}}\textrm{d}\Sigma_{\mu}J^{\mu}\,$, (104a)
$(\Delta\mathcal{E})_{\mathcal{H}}={}$ $\int_{\tilde{\mathcal{H}}}\textrm{d}\Sigma_{\mu}J^{\mu}\,$, (104b)
$(\Delta\mathcal{E})_{\chi}={}$ $\int_{\tilde{\chi}}\textrm{d}\Sigma_{\mu}J^{\mu}\,.$ (104c)
@@
$={}$ $r^{2}f\oint_{S^{2}}\textrm{d}\Omega\langle\Phi_{r}\rangle|_{r=r_{0}}\,$, (107)
where $\Phi_{r}=k^{\mu}u^{\nu}T_{\mu\nu}$, with timelike normal $u^{\mu}\partial_{\mu}=f^{-1/2}\partial_{t}$, and where $\langle\cdot\rangle$ denotes time-averaging over $\Delta t$. We notice that, by virtue of the conservation law in Eq. (106), this quantity is the same regardless of the extraction radius $r_{0}$ selected. In practice, this is the most practical way to calculate $\mathcal{F}_{\mathcal{H}}$, while its theoretical lack of $r_{0}$ dependence provides a useful consistency check for numerical solutions.
@@ Appendix D Self-consistency of decoupling approximation
The perturbative expansion applies when the Proca test field is “small” compared to the gravitational background. To define a notion of “smallness” we turn to Einstein’s equations
@@
$\epsilon\approx{}$ $\frac{r^{3}}{M}\rho$
$\approx{}$ $10^{-30}\left(\frac{\rho_{c}}{1M_{\odot}/\text{ pc}^{3}}\right)\left(\frac{M}{M_{\odot}}\right)^{2}\left(\frac{r}{r_{\rm{s}}}\right)^{3/2}\,.$ (114)